# The Domain Mixed Unit: A New Neural Arithmetic Layer

### v1

@@ Abstract
The Domain Mixed Unit (DMU) is a new neural arithmetic unit that learns a single parameter gate G that mixes a state between log-space and linear-space representations while performing either addition (DMU_add) or subtraction (DMU_sub) in said space. These are the two initializations proposed for the DMU: one covering addition and multiplication, and another covering subtraction and division. The DMU achieves state-of-the-art performance on the NALM Benchmark, a dataset designed to test the ability of neural arithmetic units to generalize arithmetic operations, specifically performing with the highest percentage solved over all seeds on multiplication and division. The DMU will be submitted as a pull request to the open-source NALM benchmark, and its code is currently available on GitHub under the forked repository: https://github.com/marict?tab=repositories.
@@ 1 Neural Arithmetic Units
Neural Arithmetic Units, introduced in [1], provide a way to encode arithmetic operations into a differentiable layer in an interpretable fashion. The idea is to place these layers inside larger models (either frozen, randomly initialized and trained, or initialized to specific operations and trained further), such as language models, to enhance their ability to perform simple arithmetic.
It has been shown that small models often struggle with arithmetic [3, 4, 5]. Research from Anthropic further indicates that large models form neural circuits capable of arithmetic, but in convoluted ways that fail to generalize with precision [6]. If given interpretable networks that can fully represent arithmetic operations to high precision, these inductive bias circuits may be used instead of the naturally occurring structures, resulting in higher performance without having to call external tools and reducing the number of tokens needed in reasoning chains.
@@ 3 Domain Mixed Unit
The proposed architecture in this paper, the Domain Mixed Unit (DMU), borrows several elements from prior work and expands upon the design. In the simplified configuration used in the evaluation on the NALM dataset, the operand selector $O$ is frozen to canonical patterns $[1,1]$ for addition and [1,1][1,-1] for subtraction, the gating parameter $G$ then determines how to mix linear- and log-domain computations for two selected inputs.
@@ 3.1 Structure
First, the unit performs the selected operation in both the linear and space on the input magnitudes using an operand selector matrix $\mathbf{O}$. This yields a linear result, $y_{\text{lin}}$, and a log-space result, $y_{\text{log}}$.
@@
$y_{\text{log}}$ $=\mathbf{O}\cdot\log(|\mathbf{x}|)$ (Log Space operation)
@@
The magnitudes are mixed in log space to prevent value blowouts and gradient explosions. To achieve this, the result of the linear path, $y_{\text{lin}}$, is first mapped to a non-negative value using a smooth absolute value approximation and then transformed into log space:
@@
$\hat{y}_{\text{lin}}$ $=\log(|y_{\text{lin}}|_{\epsilon})$ (Map linear result to log space)
The final mixed magnitude, $M_{\text{log}}$, is a weighted sum of the two paths, computed in log space:
$M_{\text{log}}=G_{\text{lin}}\cdot\hat{y}_{\text{lin}}+G_{\text{log}}\cdot y_{\text{log}}$
@@
Algorithm 1 Domain Mixed Unit (DMU) Forward Pass
1:function DMU($x,O,G,T,\epsilon$)
2:  $\triangleright$ Decompose input vector $x\in\mathbb{R}^{2}$
3:  $x_{\text{mag}}\leftarrow|x|$
4:  $x_{\text{sign}}\leftarrow\text{sign}(x)$
5:
6:  $\triangleright$ Calculate results in linear and log space for magnitudes
7:  $r_{\text{lin}}\leftarrow O\cdot x_{\text{mag}}$
8:  $r_{\text{log}}\leftarrow O\cdot\log(x_{\text{mag}})$
9:
10:  $\triangleright$ Calculate corresponding signs for both spaces
11:  $\text{sign}_{\text{lin}}\leftarrow\tanh(r_{\text{lin}}/T)$
12:  $n\leftarrow 0.5\times(1.0-x_{\text{sign}})$
13:  $m\leftarrow\sum n$
14:  $\text{sign}_{\text{log}}\leftarrow\cos(\pi\cdot m)$
15:
16:  $\triangleright$ Mix the signs using the gate G = $G_{\text{lin}}$, $G_{\text{log}}$]
17:  $V_{\text{sign}}\leftarrow G_{\text{lin}}\cdot\text{sign}_{\text{lin}}+G_{\text{log}}\cdot\text{sign}_{\text{log}}$
18:
19:  $\triangleright$ Mix the magnitudes in log space
20:  $\text{mag}_{\text{lin}}\leftarrow\sqrt{r_{\text{lin}}^{2}+\epsilon}$ $\triangleright$ Smooth absolute value
21:  $l_{\text{lin\_log}}\leftarrow\log(\text{mag}_{\text{lin}})$ $\triangleright$ Convert linear result to log space
22:  $m_{\text{log}}\leftarrow G_{\text{log}}\cdot r_{\text{log}}+G_{\text{lin}}\cdot l_{\text{lin\_log}}$
23:
24:  $\triangleright$ Convert final magnitude back to linear space
25:  $V_{\text{mag}}\leftarrow e^{m_{\text{log}}}$
26:
27:  $\triangleright$ Combine final sign and magnitude
28:  $\text{final\_value}\leftarrow V_{\text{sign}}\cdot V_{\text{mag}}$
29:
30:  return final_value
31:end function
@@ 4 Results
4.1 Addition Task Results

### v3

@@ Abstract
The Domain Mixed Unit (DMU) is a new neural arithmetic unit that learns a single parameter gate G that mixes a state between log-space and linear-space representations while performing either addition (DMU_add) or subtraction (DMU_sub) in said space. These are the two initializations proposed for the DMU: one covering addition and multiplication, and another covering subtraction and division. The DMU achieves state-of-the-art performance on the NALM Benchmark, a dataset designed to test the ability of neural arithmetic units to generalize arithmetic operations, specifically performing with the highest percentage solved over all seeds on multiplication and division. The DMU will be submitted as a pull request to the open-source NALM benchmark, and its code is currently available on GitHub under the forked repository: https://github.com/marict/nalm-benchmark.
@@ 1 Neural Arithmetic Units
Neural Arithmetic Units (NAUs) are specialized sub-units or networks designed to interpretably represent arithmetic operations while maintaining differentiability, allowing gradients to flow through them during training. These units can be integrated into larger neural architectures to provide explicit arithmetic capabilities.
2 Justification
The hypothesis underlying this work is that the inclusion of Neural Arithmetic Units (NAUs) will improve a model’s ability to perform arithmetic operations. It has been shown that small models often struggle with arithmetic [4, 5, 6]. Research from Anthropic further indicates that large models form neural circuits capable of arithmetic, but in convoluted ways that fail to generalize with precision [7].
By providing models with explicit arithmetic layers that have stronger inductive biases toward interpretable arithmetic operations, we can potentially guide the model to utilize these more principled structures instead of the naturally occurring but less reliable ones. The idea is to place these layers inside larger models (either frozen, randomly initialized and trained, or initialized to specific operations and trained further), such as language models, to enhance their ability to perform simple arithmetic.
This approach offers several advantages: (1) improved arithmetic performance without requiring external tool calls, (2) reduced token usage in reasoning chains, and (3) enhanced interpretability of the model’s arithmetic processing. If given interpretable networks that can fully represent arithmetic operations to high precision, these inductive bias circuits may be used instead of the naturally occurring structures, resulting in higher performance. The DMU represents an attempt to create such an inductive bias that can be seamlessly integrated into existing neural architectures while maintaining differentiability and end-to-end trainability.
@@ 3.5 Golden Ratio NALU (G-NALU)
4 NALM Dataset Task
The Neural Arithmetic Logic Modules (NALM) benchmark [1] is designed to evaluate the ability of neural arithmetic units to generalize arithmetic operations beyond their training distributions. The task structure is straightforward: given an input vector $\mathbf{x}\in\mathbb{R}^{2}$, the model must produce a scalar output $y\in\mathbb{R}^{1}$ where $y$ represents the result of applying a specific arithmetic operation to the components of $\mathbf{x}$.
Crucially, only one arithmetic operation is benchmarked at a time. For a given evaluation, the model is trained and tested exclusively on one of the four fundamental operations: addition ($x_{1}+x_{2}$), subtraction ($x_{1}-x_{2}$), multiplication ($x_{1}\times x_{2}$), or division ($x_{1}\div x_{2}$). This allows for focused evaluation of how well each neural arithmetic unit can learn and generalize the specific mathematical relationship.
Generalization performance is evaluated using a weight perturbation method where optimal solution weights are perturbed by $\epsilon=1\times 10^{-5}$, and extrapolation loss is computed over 1 million examples, resulting in a specific threshold for each neural arithmetic unit, range, and operation combination. The threshold values used for the DMU evaluation are provided in the appendix, with representative values ranging from $10^{-9}$ for the most precise operations to $10^{-1}$ for the most challenging extrapolation scenarios.
$\boxed{\begin{array}[]{c}\mathbf{x}\in\mathbb{R}^{2}\\ \text{(Input)}\end{array}}\xrightarrow{\text{Operation}}\boxed{\begin{array}[]{c}o\in\{+,-,\times,\div\}\\ \text{(Function)}\end{array}}\xrightarrow{\text{Apply}}\boxed{\begin{array}[]{c}y\in\mathbb{R}^{1}\\ \text{(Ground Truth)}\end{array}}$
Examples:
$[2.5,1.3]$ $\xrightarrow{\text{Addition}}3.8$
$[2.5,1.3]$ $\xrightarrow{\text{Multiplication}}3.25$
Figure 1: NALM Dataset Task Structure: The task maps a 2D input vector to a 1D output through a single arithmetic operation. Only one operation is evaluated per experiment.
@@ 5 Domain Mixed Unit
The proposed architecture in this paper, the Domain Mixed Unit (DMU), borrows several elements from prior work and expands upon the design.
@@ 5.1 Structure
First, the unit performs the selected operation in both the linear and space on the input magnitudes using an operand selector matrix $\mathbf{O}$. This yields a linear result, $y_{\text{lin}}$, and a log-space result, $y_{\text{log}}$. The log-space operation includes clamping to prevent numerical instabilities.
@@
$y_{\text{log}}$ $=\mathbf{O}\cdot\log(\text{clamp}(|\mathbf{x}|,\text{min}=\epsilon_{\text{mag}}))$ (Log Space operation)
@@
The magnitudes are mixed in log space to prevent value blowouts and gradient explosions. To achieve this, the result of the linear path, $y_{\text{lin}}$, is first mapped to a non-negative value using a smooth absolute value approximation and then transformed into log space with clamping:
@@
$\hat{y}_{\text{lin}}$ $=\log(\text{clamp}(|y_{\text{lin}}|_{\epsilon},\text{min}=\epsilon_{\text{mag}}))$ (Map linear result to log space)
$\tilde{y}_{\text{log}}$ $=\text{clamp}(y_{\text{log}},\text{min}=-L,\text{max}=L)$ (Clamp log values)
The final mixed magnitude, $M_{\text{log}}$, is a weighted sum of the two paths, computed in log space with final clamping:
$M_{\text{log}}^{\prime}$ $=G_{\text{lin}}\cdot\hat{y}_{\text{lin}}+G_{\text{log}}\cdot\tilde{y}_{\text{log}}$
$M_{\text{log}}$ $=\text{clamp}(M_{\text{log}}^{\prime},\text{min}=-L,\text{max}=L)$ (Final magnitude clamp)
@@
Algorithm: Domain Mixed Unit (DMU) Forward Pass
Function DMU(x, O, G, T, epsilon, mag_min, log_lim):
    // Decompose input vector x in R^2
    x_mag = |x|
    x_sign = sign(x)
    // Calculate results in linear and log space for magnitudes
    r_lin = O * x_mag
    r_log = O * log(clamp(x_mag, min=mag_min))  // Clamp before log
    // Calculate corresponding signs for both spaces
    sign_lin = tanh(r_lin / T)
    n = 0.5 * (1.0 - x_sign)
    m = sumn
    sign_log = cos(pi * m)
    // Mix the signs using the gate G = [G_lin, G_log]
    V_sign = G_lin * sign_lin + G_log * sign_log
    // Mix the magnitudes in log space
    mag_lin = sqrt(r_lin^2 + epsilon)  // Smooth absolute value
    l_lin = log(clamp(mag_lin, min=mag_min))  // Clamp before log
    l_log = clamp(r_log, min=-log_lim, max=log_lim)  // Clamp log values
    m_log = G_log * l_log + G_lin * l_lin
    m_log = clamp(m_log, min=-log_lim, max=log_lim)  // Final clamp
    // Convert final magnitude back to linear space
    V_mag = e^(m_log)
    // Combine final sign and magnitude
    final_value = V_sign * V_mag
    Return final_value
5.2 NALM Configuration
In the simplified configuration used for evaluation on the NALM dataset, the operand selector $O\in\{1,-1\}^{2}$ is frozen to canonical patterns $[1,1]$ for addition (DMU_add) and [1,1][1,-1] for subtraction (DMU_sub). With these fixed operand patterns, the gating parameter $G$ becomes the primary learnable component that determines how to mix linear- and log-domain computations for the two selected inputs.
@@ 6 Results
6.1 Statistical Analysis
Comprehensive evaluation across 900 total experiments (25 seeds per 4 operations and 9 ranges) reveals the DMU’s exceptional performance consistency with 100% success rate across all operations.
Table 2: DMU Performance Analysis
Operation Mean Conv. Time Mean Sparsity Mean Extrapolation Error
ADD 2170 0.281 $3.3\times 10^{-6}$
SUB 7842 0.280 $1.5\times 10^{-6}$
MUL 5626 0.266 $1.4\times 10^{-5}$
DIV 1400 0.277 $7.3\times 10^{-8}$
The mean sparsity values indicate effective parameter utilization, while the extremely low extrapolation error values demonstrate precise convergence to target functions across all evaluation ranges.
6.2 Operation Task Results
@@
7 Loss Landscapes
The exceptional performance of DMU_add and DMU_sub can be attributed to their operation in an optimal parameter space. When the operand selector $O$ is frozen to the correct canonical patterns and only the gating parameter $G$ is learned, the loss landscape forms a smooth, convex function that enables reliable convergence to the global optimum.
However, when both $O$ and $G$ are simultaneously optimized, the loss landscape becomes significantly more complex, characterized by numerous saddle points that impede convergence. This phenomenon is particularly pronounced in subtraction tasks, where the loss surface exhibits slopes that push the model parameters toward incorrect values of $O[1]$, as illustrated in Figure 2.
Figure 2: Loss landscape for subtraction operation showing a problematic slope that pushes the unit toward an incorrect $O[1]$ value.
The presence of these saddle points reveals a fundamental challenge in designing universal neural arithmetic units. While the DMU achieves perfect performance on individual operations through careful parameter initialization and selective optimization, creating a single unit capable of learning all four arithmetic operations simultaneously requires sophisticated methods to smooth and resolve these problematic regions in the loss landscape.
If a Neural Arithmetic Unit is to be created that performs well on all operations without operation-specific initialization, it must incorporate mechanisms to navigate around or eliminate these saddle points, potentially through architectural innovations, specialized training procedures, or hybrid approaches that combine the benefits of both fixed and learnable components.
@@ 9 Extensibility
10 Future Work
Two key research directions emerge from this work:
  1. 1.
    Integration with Large Language Models: Document whether LLM performance on mathematics is improved by the inclusion of NAUs. This can be achieved by analyzing pretraining performance on math-heavy datasets like Proof-Pile-2 [8], and the interpretability of NAUs should allow us to observe the operations being encoded from their respective input contexts.
  2. 2.
    Unified Arithmetic Operations: Resolve the saddle point issues in internal operation space to allow a single NAU to extrapolate on all arithmetic operations, eliminating the need for operation-specific configurations like DMU_add and DMU_sub.
Appendix
DMU Convergence Thresholds
The following table shows the convergence thresholds used to evaluate DMU performance across different operations and ranges. These thresholds were determined using the weight perturbation method described in Section 4.
Table 7: DMU Convergence Thresholds by Operation and Range
Range Addition Subtraction Multiplication Division
sym 7.55e-07 1.31e-07 1.27e-05 4.55e-08
neg 1.14e-06 9.44e-08 2.35e-05 6.59e-08
pos 3.68e-07 1.67e-07 2.04e-06 2.53e-08
n10 1.13e-06 1.76e-07 1.83e-05 9.25e-08
p01 2.61e-08 3.42e-08 4.31e-09 1.06e-07
n01 3.64e-07 7.68e-09 6.13e-08 4.07e-07
p11 2.63e-07 3.04e-07 1.39e-06 2.26e-08
n20 1.96e-05 8.12e-06 1.99e-01 1.99e-08
p20 2.73e-04 9.47e-06 6.67e-02 3.19e-08
Mean 3.30e-05 2.06e-06 2.95e-02 9.07e-08